### v1

# Self-Interacting Dark Matter with Mass Segregation: A Unified Explanation of Dwarf Cores and Small-Scale Lenses

Abstract
In two-component SIDM models with inter-species interactions, mass segregation arises naturally from collisional relaxation, enhancing central densities and gravothermal evolution without requiring large cross sections. We propose a model with velocity-dependent interactions, both within and between species, that connects observations across several halo mass scales while remaining consistent with cluster-scale constraints. This combination enables modest mass segregation in low-mass and typical-concentration halos, consistent with recent dwarf galaxy clustering measurements. Using cosmological zoom-in simulations and controlled isolated halo studies, we show that this model produces dwarf galaxy cores that grow over time, explains the structure of dark perturbers observed in strong lensing systems, and significantly increases the number and efficiency of small-scale lenses, consistent with the galaxy-galaxy strong lensing excess reported in clusters. Our results establish mass segregation in two-component SIDM as a self-consistent and testable model capable of simultaneously addressing multiple small-scale challenges in structure formation.
Introduction— Observations of large-scale structure have established the standard cosmological model with cold dark matter (CDM). However, on small scales, persistent discrepancies continue to challenge this framework, motivating alternative interpretations Bullock and Boylan-Kolchin (2017); Salucci (2019); Gentile et al. (2004); van Dokkum et al. (2022); Bonaca et al. (2018). Among these, self-interacting dark matter (SIDM) has long been considered a promising candidate Spergel and Steinhardt (2000); Tulin and Yu (2018). Through gravothermal evolution, SIDM can naturally produce both cored and cuspy halo density profiles, providing a flexible mechanism to explain the diversity of galactic rotation curves, the extreme central densities observed in ultra-diffuse and ultra-compact galaxies, and the early formation of supermassive black holes Kaplinghat et al. (2016); Robertson et al. (2019); Despali et al. (2019); Essig et al. (2019); Santos-Santos et al. (2020); Turner et al. (2021); Fischer et al. (2021); Slone et al. (2023); Zeng et al. (2022); Gilman et al. (2021); Correa et al. (2022); Yang and Yu (2022); Adhikari et al. (2022); Outmezguine et al. (2023); Silverman et al. (2022); Yang et al. (2024); Zhong et al. (2023); Zeng et al. (2025); Zhang et al. (2025a, 2024); Ragagnin et al. (2024); Nadler et al. (2025); Kong and Yu (2025); Jiang et al. (2025). It has also been proposed as a solution to the final-parsec problem in binary black hole mergers Alonso-Álvarez et al. (2024). Recently, a striking development by Zhang et al. Zhang et al. (2025b) revealed an unexpected clustering pattern in dwarf galaxies, favoring a self-interaction cross section of $\sigma/m\sim 0.3~{}\rm cm^{2}/g$ in their host halos.
Despite accumulating evidence, a unified SIDM explanation for all small-scale phenomena remains elusive. In particular, the relatively small cross section inferred from dwarf galaxy clustering appears inconsistent with the onset of core collapse required to explain dense subhalos observed via strong lensing, which demands $\sigma/m\gtrsim 100~{}\rm cm^{2}/g$ Nadler et al. (2023); Li et al. (2025). Such compact substructures have been inferred in systems like SDSS J0946+1006 and JVAS B1938+666  Vegetti et al. (2010, 2012); Despali et al. (2024); Ballard et al. (2024); Enzi et al. (2024). Moreover, core collapse alone cannot fully account for the elevated abundance of small-scale lenses reported in galaxy-galaxy strong lensing (GGSL) by Meneghetti et al Meneghetti et al. (2020, 2022, 2023), which exceeds CDM expectations by an order of magnitude. Ref. Dutra et al. (2025) shows that resolving this tension requires extreme central densities in all subhalos, in conflict with the cored profiles needed in dwarfs. Ref. Yang and Yu (2021) further finds that SIDM halos with baryons primarily enhance two-image lensing events, rather than the four-image ones observed.
In this work, we demonstrate that SIDM with mass segregation provides a natural unified framework for resolving these tensions. Mass segregation in multi-component SIDM models arise as a consequence of collisions among particles of different masses. In halos containing two dark matter species with inter-species scattering, energy transfer drives the heavier component toward the center, enhancing central densities Yang et al. (2025a); Patil and Fischer (2025). This process is negligible in CDM due to inefficient gravitational relaxation among particles (unlike stars Zhong and Shapiro (2025)), but becomes significant in SIDM, where collisional interactions dominate. Previous simulations at Milky Way scales Yang et al. (2025a) demonstrated that this mechanism reshapes inner halo profiles under strong scattering. Here, we extend this framework to cosmological simulations of galaxy clusters and controlled simulations of dwarf-scale halos. We show that even modest inter-species interactions can induce mass segregation in massive or high-concentration halos, increasing central densities while leaving dwarf halos largely unaffected. Crucially, unlike in one-component SIDM, the mass enclosed within the Einstein radius can also grow, amplifying strong lensing signals. Our results point to a unified scenario in which both cored dwarf galaxies and cuspy cluster substructures emerge naturally from SIDM with mass segregation.
Figure 1: Effective cross section per mass $\sigma_{\rm eff}/m$ as a function of $V_{\rm max}$ for various SIDM models. The self-interaction among the heavier component ($\chi_{H}-\chi_{H}$) and the inter-species interaction ($\chi_{H}-\chi_{L}$) in the SIDM2v model are represented by red and magenta solid curves. This model is featured by its capability of reproducing SIDM effects analogous to one dark matter component with $\sigma/m=0.3~{}{\rm cm^{2}/g}$ (SIDM03) in massive dwarf halos of mass around $M_{h}=10^{11}~{}M_{\odot}$, as favored by the findings of Ref. Zhang et al. (2025b). The two orange stars feature the two kinds of strengths in the two-component SIDM2c model, which takes constant effective cross sections as in SIDM2v in a $M_{h}=10^{11}~{}\rm M_{\odot}$ halo with median concentration. For $V_{\rm max}\gtrsim 200~{}\mathrm{km/s}$, the inter-species interaction dominates over the intra-species interactions and remains significant up to the cluster scale at $V_{\rm max}\approx 1600~{}\mathrm{km/s}$. In clusters, the coexistence of both types of scattering yields density profiles consistent with observations. Particularly, the intra-species interaction remains well below $0.1~{}\rm cm^{2}/g$, as indicated by the light-red arrow.
A two-component SIDM model with mass segregation— To incorporate mass segregation, we consider a two-component SIDM model in which the heavier species, $\chi_{H}$, has mass $m_{H}=3m_{L}$, with $\chi_{L}$ denoting the lighter component. We assume equal number densities and identical initial spatial distributions for both species. We introduce a model, SIDM2v, featuring velocity-dependent interactions for both identical (intra-species) and distinguishable (inter-species) scatterings. While both cross sections decrease with velocity, the inter-species interaction dominates at high velocities ($V_{\rm max}\gtrsim 200~{}\mathrm{km/s}$). Specifically, we adopt the Møller and Rutherford parametrizations for identical and distinguishable particle scatterings, following equations in Refs. Yang et al. (2025a); Yang and Yu (2022). The intra-species interaction is characterized by $\sigma_{0}/m_{H}=6.89~{}\rm cm^{2}/g$ and a velocity scale $w=275~{}\rm km/s$ (denoted $w_{1}$), while the inter-species interaction uses $\sigma_{0}/m_{H}=1.125~{}\rm cm^{2}/g$ and $w=2200~{}\rm km/s$. We assume that the same particle mediates all intra-species interactions, so that the $\chi_{L}-\chi_{L}$ interaction shares the same $\sigma_{0}$, with a velocity scale $w=(m_{H}/m_{L})w_{1}$.
Figure 1 illustrates our model in terms of effective cross sections, defined as velocity- and angle-averaged differential cross sections weighted by a $v^{5}\sin^{2}\theta$ kernel and convolved with the dark matter velocity distribution. The $x$-axis corresponds to halo velocity scales via their maximum circular velocity, $V_{\rm max}$. The two orange stars indicate the intra- and inter-species interactions in a SIDM2c model without velocity dependence. Their strengths are set by matching the effective cross sections of SIDM2v in a benchmark halo of mass $10^{11}~{}\rm M_{\odot}$, yielding $\sigma_{0}/m=2.25~{}\rm cm^{2}/g$ for intra-species and $\sigma_{0}/m=1.125~{}\rm cm^{2}/g$ for inter-species interactions. This setup leads to time-dependent core formation in most dwarf-hosting halos of mass $\sim 10^{11}~{}\rm M_{\odot}$, considering a single-component SIDM03 model with $\sigma/m=0.3~{}\rm cm^{2}/g$, consistent with the prediction of Ref. Zhang et al. (2025b) and shown as a dashed black line here.
To examine small-scale effects in detail, we perform high-resolution simulations of isolated dwarf halos in SIDM2c, which replicate the relevant interaction regime in SIDM2v. These complement our cosmological simulations near their resolution limit. At higher masses ($M\gtrsim 10^{12}~{}\mathrm{M}_{\odot}$), the intra-species cross section (red curve) drops more steeply than the inter-species one (magenta), leading to enhanced central densities. We also consider a model, SIDMx, that includes only inter-species interactions from SIDM2c, representing an extreme scenario of mass segregation. Notably, both intra- and inter-species interactions are weak yet effective under SIDM2v in clusters. Inter-species interactions drive condensation, which competes with core formation induced by intra-species self-interactions, resulting in a density profile consistent with the $\sigma/m<0.1~{}\rm cm^{2}/g$ constraint from cluster lensing observations Andrade et al. (2021); Kaplinghat et al. (2016); Andrade et al. (2019). We provide an additional supportive test in the Supplemental Material.
Our simulations are based on the SIDM module developed in Refs. Yang and Yu (2022); Yang et al. (2023), extended to handle two dark matter species with unequal masses. This module has been validated in both isolated and cosmological contexts. For isolated halos, we generate NFW profiles in hydrostatic equilibrium using SpherIC Garrison-Kimmel et al. (2013), simulating with $10^{6}$ particles per run. For cosmological simulations, we initialize a periodic box of size $400~{}{\rm Mpc}/h$, and perform zoom-in simulations of a $20~{}{\rm Mpc}/h$ region centered on a host halo of mass $10^{15}~{}{\rm M_{\odot}}/h$, where $h$ is the dimensionless Hubble parameter. The highest-resolution particles have average mass $m_{0}\equiv(m_{1}+m_{2})/2\approx 7\times 10^{7}~{}{\rm M_{\odot}}/h$. As shown in Ref. Yang et al. (2025a), mass segregation is negligible in CDM when particle masses are realistic, so we simulate CDM using a single-component model. For one-component SIDM in isolated halos, we employ the parametric framework of Ref. Yang et al. (2024), which has been extensively tested across controlled and cosmological settings Yang et al. (2025b).
Figure 2: Density profiles of simulated $10^{11}~{}\rm M_{\odot}$ halos with median and ±$\pm 1\sigma$ concentrations at $t=5~{}{\rm Gyr}$ under SIDM03 (dashed) and SIDM2c (solid). The closeness between dashed and solid curves demonstrates that two-component SIDM models retain core formation comparable to the one-component case, consistent with dwarf galaxy clustering observations. The inset further shows that the core formation is time-dependent in two-component models. In the $+1\sigma$ concentration case, we observe that the central densities have clearly increased past their minimum, suggesting an accelerated gravothermal evolution in high concentration halos.
Dwarf scale cores— Ref. Zhang et al. (2025b) finds that diffuse dwarf galaxies exhibit unexpectedly stronger large-scale clustering than denser ones. This can be naturally explained in SIDM, where halos in the core-forming phase develop larger cores if they evolve longer. Since halos in overdense regions form earlier, they are more likely to host diffuse galaxies, leading to enhanced clustering. To demonstrate that our proposed two-component models can reproduce this behavior, we simulate isolated halos with mass $M=10^{11}~{}\rm M_{\odot}$, which host the dwarfs in Ref. Zhang et al. (2025b), using SIDM2c. These simulations show that the halos develop time-dependent cores, similar to those in the one-component model with $\sigma/m=0.3~{}\rm cm^{2}/g$ (SIDM03), as favored in that work.
Figure 2 compares the resulting density profiles at $t=5~{}$Gyr for halos in SIDM03 (dashed), and SIDM2c (solid), with median (green), $+1\sigma$ (orange), and $-1\sigma$ (blue) concentrations, based on the redshift zero $c_{200}-M$ relation from Ref. Dutton and Macciò (2014). We find that both the one- and two-component models produce similarly sized cores, confirming that SIDM2c, and by extension its velocity-dependent counterpart SIDM2v, retains the ability to explain dwarf galaxy clustering despite its richer internal structure.
[…]
The inset panel in Fig. 2 tracks the time evolution of central densities. We observe that more concentrated halos evolve more rapidly in the two-component case, reaching the minimum central density at earlier times compared to their one-component counterparts. This suggests that mass segregation enhances the efficiency of gravothermal evolution in dense systems, driving the most extreme cases into core collapse within a Hubble time. This has important implications for the origin of the highly compact dark substructures inferred from strong lensing Vegetti and Koopmans (2009); Vegetti et al. (2010, 2012); Hezaveh et al. (2016); Minor et al. (2021); Tajalli et al. (2025). A small population of high-concentration dwarf halos can evolve into dense, collapsing configurations even with modest self-interaction cross sections that are insufficient to induce core collapse in one-component SIDM, as explored below.
Figure 4: Comparison of simulated gravitational lenses in CDM (left), SIDM2v (middle), and SIDMx (right). Top panels show tangential (opaque) and radial (40% transparent) critical curves overlaid on projected density maps; bottom panels show corresponding caustics. The two-component models produce significantly more secondary caustics due to dense substructures. The GGSL cross section ($\sigma_{\rm GGSL}$) is computed summing over the areas enclosed by the secondary caustics whose effective Einstein radii are larger than $0.5^{\prime\prime}$ and residing within the $(-100^{\prime\prime},100^{\prime\prime})$ square region centered on the host halo. The SIDM2v has $\sigma_{\rm GGSL}=6.76$, being about $13$ times that of the CDM case and aligns with the observation by Ref. Meneghetti et al. (2020). The SIDMx model represents an extreme case with only inter-species interaction and $\sigma_{\rm GGSL}=12.8$.
Strong lensing perturber— SDSSJ0946+1006 provides a pronounced example of a compact, dark substructure revealed through strong lensing, with image reconstruction indicating a steep projected density slope, $\gamma_{\rm 2D}\sim-1.25$, and high mass within the inner kiloparsec Minor et al. (2021). Such subhalos challenge both CDM and single-component SIDM models. In CDM, it is extremely difficult to generate sufficiently dense and compact subhalos without invoking rare statistical fluctuations. While core-collapsing SIDM models can, in principle, produce such dense objects, prior studies have shown that cross sections of $\gtrsim 100~{}\rm cm^{2}/g$ are required for a $10^{11}\,M_{\odot}$ halo to reach collapse Nadler et al. (2023); Li et al. (2025), creating tension with the much smaller cross sections inferred from dwarf galaxy clustering. Moreover, halos that reach such densities would typically be expected to host observable galaxies, in conflict with the non-detection of luminous counterparts.
To illustrate how SIDM with mass segregation can alleviate these tensions, we simulate three benchmark halos under SIDM2c with masses $M=5\times 10^{10}$, $8\times 10^{10}$, and $10^{11}~{}\rm M_{\odot}$, labeled M05, M08, and M11, respectively. Each is initialized with a concentration $2.5\sigma$ above the cosmological median.
Figure 3 shows the evolution of the benchmark halos in the plane of projected inner mass (within 1 kpc) versus average projected slope $\gamma_{\rm 2D}$, with arrows indicating their trajectories over time. All three evolve toward the region favored by the SDSSJ0946+1006 perturber, with M08 passing through the 95% confidence contour at $t\approx 7.4$ Gyr. This demonstrates that mass segregation in two-component SIDM can drive certain halos into dense, lensing-relevant configurations even at modest cross sections. By contrast, in SIDM03, the same halos remain in the core-forming phase and evolve away from the dense regime required to match the perturber. To connect with the cosmological context, we overlay subhalos from the SIDM2v simulation. Despite resolution limits, these subhalos extend to lower slopes and higher inner masses, with several candidates residing within the region favored by the SDSSJ0946+1006 perturber, supporting the viability of forming compact, dark structures in SIDM models with mass segregation.
Small-Scale Lenses and Lensing Cross Sections— The capability of enhancing inner masses in two-component models also makes subhalos more efficient lenses and addresses the challenge reported in Ref. Meneghetti et al. (2020), where the GGSL cross section, defined as the area enclosed by tangential caustics induced by subhalos, is an order of magnitude larger in observations. This challenge remains difficult to reconcile with current models. Simulations with baryonic physics can increase subhalo densities, but the effect is often driven by one or two massive satellites and fails to robustly match observed statistics Bahé (2021); Robertson (2021); Dutra et al. (2025). In SIDM, the core collapse has demonstrated potential, but achieving it across all subhalos with inner density slopes steeper than $r^{-2.5}$ remains challenging Dutra et al. (2025). Moreover, when baryons are included, core-collapsed subhalos tend to enhance radial caustics rather than the tangential ones relevant for four-image systems Yang and Yu (2021). In contrast, mass segregation naturally increases the central densities and steepens the inner density slopes of subhalos. To facilitate comparison with the GGSL measurment in Ref. Meneghetti et al. (2020), we fix the redshift of our host system at $z=0.439$ and compute lensing properties by including all subhalos within a $5~{}{\rm Mpc}/h$ sphere. The lensing potential and deflection angles are primarily modeled analytically using the parametric SIDM halo profile developed in Hou et al. (2025); Yang et al. (2024), which allows for efficient calculation of lensing observables across a wide range of halo masses and internal structures. See the Supplemental Material for details.
[…]
Conclusion— Our simulations show that SIDM with mass segregation retains cores in most dwarf halos, accelerates inner condensation in high-concentration halos, and yields denser subhalos in clusters. These results highlight SIDM2v as a unified and testable framework for resolving multiple small-scale challenges. Looking ahead, our framework may also shed light on the formation of early massive galaxies and supermassive black holes, including the so-called “little red dots”, as mass segregation remains effective at high redshifts. Our results suggest that small-scale structure may already be revealing not only the strength but also the multi-component nature of dark matter.
Acknowledgements.
We thank Moritz Fischer, Guoliang Li, Nan Li, Ran Li, Shihong Liao, Priyamvada Natarajan, Huiyuan Wang, Hai-Bo Yu, and Yi-Ming Zhong for helpful discussion. The authors were supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), Chinese Academy of Sciences (CAS), and the Project for Young Scientists in Basic Research of CAS (No. YSBR-092).
Appendix A SUPPLEMENTAL MATERIAL
Appendix B Cluster density profiles and constraint
Figure 5: Left: Density profiles of cluster host halos in the CDM, SIDM2v, and SIDMx simulations (dashed). All profiles are well fitted by a two-component parametric SIDM model (solid) Hou et al. (2025); Yang et al. (2024). Right: Evolution of an isolated cluster host halo, taking scale parameters from the cluster halo in the CDM simulation, under SIDM2v. At either $t=5~{}\rm Gyr$ (solid) or $14~{}\rm Gyr$ (dotted), the resulting core (blue) is much smaller than that produced in a one-component model with $\sigma/m=0.1~{}\rm cm^{2}/g$ (SIDM01, red), demonstrating that SIDM2v remains consistent with cluster-scale constraint.
The reconstruction of cluster mass profiles has presented one of the most robust constraints on SIDM in clusters, which varies but is generically around $0.1~{}\rm cm^{2}/g$ Andrade et al. (2021); Kaplinghat et al. (2016); Andrade et al. (2019). Core formation in clusters is efficient and only moderately influenced by baryon-induced contraction. For a constant, small self-interaction cross section per unit mass, the rate of gravothermal evolution, quantified by the inverse of the core-collapse time, is proportional to $\rho_{s}^{3/2}r_{s}$, implying that more massive halos evolve more rapidly. Moreover, the stellar mass in clusters generally constitutes only about $0.05\%$ of the total halo mass, and thus contributes negligibly to contraction.
The presence of mass segregation introduces a mechanism that enhances inner densities, counteracting core formation and thereby alleviating the constraint. In SIDM2v, for example, the inter-species interaction, represented by the magenta curve in Fig. 1 around $V_{\rm max}\approx 1600~{}\rm km/s$, exceeds $0.1~{}\rm cm^{2}/g$ and is significantly stronger than the intra-species interaction. This allows it to compete with the efficient core formation.
To demonstrate the interplay between inter- and intra-species interactions, and to show that cluster halos in SIDM2v remain consistent with the current $0.1~{}\rm cm^{2}/g$ constraint, we present the density profiles of host halos from cosmological simulations in the left panel of Fig. 5. We also simulate an isolated cluster halo under SIDM2v, using the scale parameters from the CDM cluster halo as a reference. The resulting density profiles are denser than those of the one-component SIDM01 model at both $t=5$ and $14~{}\rm Gyr$. Since the observational constraint is based on the inner mass distribution, our results indicate that clusters in SIDM2v are less constrained than in SIDM01.
Appendix C Calculation of strong lensing effects
Figure 6: Density profiles of all cluster subhalos with $M>10^{11}\,M_{\odot}/h$ in the CDM, SIDM2v, and SIDMx simulations (dashed). Solid curves show corresponding fits using the parametric SIDM model of Refs. Hou et al. (2025); Yang et al. (2024). Subpanels display the relative deviation between simulation and fit, defined as $2\,(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. The parametric model accurately captures the density structure of simulated subhalos across a broad mass range.
We compute strong lensing effects using the parametric approach developed in Refs. Hou et al. (2025); Yang et al. (2024). Starting from the simulated particle data, we extract spherically averaged halo density profiles under the assumption of spherical symmetry. These profiles are then modeled using a sum of two parametric SIDM (PSIDM) halo components, each described by an analytic fitting function. For each PSIDM component, the density profile is given by
$\rho_{\rm PSIDM}(r|\rho_{s,0},r_{s,0},\tau)$ (1)
$=\frac{\rho_{s}}{\left[\left(\frac{r}{r_{s}}\right)^{4}+\left(\frac{r_{c}}{r_{s}}\right)^{4}\right]^{\frac{\gamma}{4}}\cdot\left[1+\left(\frac{r}{r_{s}}\right)^{\beta}\right]^{\frac{3-\gamma}{\beta}}},$
where $\rho_{s}=\rho_{s}(\rho_{s,0},\tau)$, $r_{s}=r_{s}(r_{s,0},\tau)$, $r_{c}=r_{c}(r_{s,0},\tau)$, $\beta=\beta(\tau)$ and $\gamma=\gamma(\tau)$ are functions of the phase parameter $\tau\in(0,1.08)$, defined explicitly in Ref. Hou et al. (2025). Since each PSIDM component is determined by three parameters, the sum of two such components involves six fitting parameters in total.
Figure 6 shows the fitting performance for cluster subhalos in three cosmological simulations: from left to right, CDM (blue), SIDM2v (red), and SIDMx (magenta). The lower panels illustrate the relative difference between the simulated (Sim) and fitted (Fit) profiles, computed as $2(\mathrm{Fit}-\mathrm{Sim})/(\mathrm{Fit}+\mathrm{Sim})$. The two-component PSIDM model reproduces the simulated profiles well, and the remaining scatter is primarily due to statistical fluctuations in the simulation data.
Based on the fitted profiles, Ref. Hou et al. (2025) provides analytic expressions for the surface density, lensing potential, and deflection angles of each halo. The linearity of the lensing equation allows us to stack their contributions for more comprehensive lensing predictions.
For halos other than the main cluster hosts, their internal ellipticities contribute subdominantly to the lensing shear compared to the external shear from the host halo. However, the mass and ellipticity of the host halo significantly affect the lensing signal and must be incorporated. We model this effect by applying a coordinate transformation that introduces a fixed ellipticity into the projected mass distribution. The ellipticity of each halo is measured by projecting the associated particles onto the lens plane, taken to be the $x$$y$ plane of the simulation box. We define the axis ratio $q=b/a$ of the projected mass distribution and compute the elliptical radial coordinate as Kormann et al. (1994)
$s=\sqrt{q^{2}(x\cos\phi+y\sin\phi)^{2}+(-x\sin\phi+y\cos\phi)^{2}},$
where $\phi$ is the angle between the $x$-axis and the semi-major axis of the lens. Substituting $s$ for $r$ in the circular lens model yields an elliptical lens, preserving the enclosed mass within isodensity contours. We apply a fast Fourier transform to the elliptical surface density to compute the deflection angle maps for the host halos. For the CDM, SIDM2v, and SIDMx clusters, the measured axis ratios are $q=0.39$, $0.42$, and $0.47$, respectively. These large ellipticities enhance the GGSL effects associated with the host halos.
Appendix D Strong lensing effects under identical host lens
Figure 7: Left: Tangential (opaque) and radial (40% transparent) critical curves (red) overlaid on projected density maps in the SIDM2v (top) and SIDMx (bottom) simulations, where the host halo has been replaced with the CDM host halo to isolate subhalo contributions. Right: Tangential caustics (green) correspond to the critical curves on the left panels. The enhancement in the number and strength of small-scale lenses remains clearly demonstrated, confirming that subhalo structure is the primary driver of GGSL cross section differences between CDM and two-component SIDM models. See text for details.
To address the excess in the GGSL cross section, we primarily focus on the influence of cluster subhalos. To assess the impact of uncertainties in modeling host halos, we perform an extreme test by adopting the CDM cluster as the host lens in all cases. This approach helps to marginalize the contribution from host halo variations.
Figure 7 shows the critical curves (left) and caustics (right) for the SIDM2v and SIDMx simulations, with their host halos replaced by the CDM host. The presence of a significant number of small-scale lenses remains visually apparent through the enclosed curves. In the SIDM2v case, more structures are visible in the upper-left region, which can be attributed to a broader spatial distribution of massive subhalos compared to SIDMx. We leave a detailed investigation of the origin of these differences, likely related to variations in accretion history, for future work with higher-resolution simulations.
Under this setup, the resulting GGSL cross sections are $0.515$, $11.7$, and $6.50$ for the CDM, SIDM2v, and SIDMx simulations, respectively. The corresponding numbers of secondary lenses are $5$, $25$, and $26$. Comparing these results with those presented in the main text, we find that the choice of cluster halo can indeed introduce substantial differences. The differences in the simulated host halos can be physical. However, future simulations with higher resolution and full baryonic physics may further refine these details. Despite this, our conclusions remain robust: the ratio of $\sigma_{\rm GGSL}$ between SIDM2v and CDM increases from $13$ to $23$. The number of small-scale lenses decreases slightly, from $30$ to $25$ in SIDM2v, and from $37$ to $26$ in SIDMx, but both remain significantly higher than the $5$ found in CDM.
Appendix E Candidates contributing to the GGSL cross sections
Figure 8: Bound subhalo mass ($M_{\rm sub}$) versus $r_{\rm max}$ for subhalos associated with the selected secondary lenses in the CDM (left), SIDM2v (middle), and SIDMx (right) simulations. Points are color-coded by the effective Einstein radius, defined as $\theta_{E}=\sqrt{A_{E}/\pi}$, where $A_{E}$ is the area enclosed by the corresponding critical curve. The subhalo mass range spans $10^{10}-10^{12}~{}\rm M_{\odot}/h$, and the resulting Einstein radii lie within $0.5^{\prime\prime}$$3^{\prime\prime}$, consistent with the findings of Ref. Dutra et al. (2025).
Ref. Dutra et al. (2025) demonstrates that the subhalos contributing to GGSL predominantly have bound masses in the range $10^{10}-10^{12}~{}\rm M_{\odot}/h$ and effective Einstein radii between $0.5^{\prime\prime}$ and $3^{\prime\prime}$, where the Einstein radius is defined as $\theta_{E}=\sqrt{A_{E}/\pi}$, with $A_{E}$ denoting the area enclosed by the critical curves. To show that the properties of the small-scale lenses in our study are consistent with these findings, we identify subhalos with effective $\theta_{E}>0.5^{\prime\prime}$ and plot their bound mass ($M_{\rm sub}$) against the radius at which their circular velocity peaks ($r_{\rm max}$) in Fig. 8, for the CDM (left), SIDM2v (middle), and SIDMx (right) simulations. The effective Einstein radii are indicated by the color scale. This result further supports the interpretation that SIDM models with mass segregation can produce subhalos consistent with GGSL observations.